OpenStudy (anonymous):

can someone check my answer? Evaluate the expression, using the given value of the variable. 4 – 2x + 5x when x = 4 A. –24 B. –18 C. 16 D. 28 I got A

TheSmartOne (thesmartone):

mhmmm

TheSmartOne (thesmartone):

show us your work :)

OpenStudy (anonymous):

okay! 4-(2x4)+(5x4)= 8+20=28 4-28= -24

OpenStudy (zenmo):

Nope

TheSmartOne (thesmartone):

Ah, you see 4 - 8 + 20 we can't do 8 + 20 we have to do - 8 + 20

TheSmartOne (thesmartone):

That's why it's best to do 4 - 8 first :)

TheSmartOne (thesmartone):

so what is 4 - 8 = ?

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

4-8= -4

TheSmartOne (thesmartone):

and -4 + 20 = ?

OpenStudy (anonymous):

16

TheSmartOne (thesmartone):

And that's your answer :)

OpenStudy (anonymous):

ohhh. thanks!!

TheSmartOne (thesmartone):

Anytime! :D

OpenStudy (zenmo):

PEMDAS: 1)Parenthesis 2)Exponents 3)Multiply 4)Divide 5)Add 6)Subtract You would do that in the order for solving these types of equations

OpenStudy (zenmo):

Whenever you're left with add and subtraction, you solve from left to right